Find the slope of the line through the following pairs of points.
(-5, -3) and (-5, 2)

Answer by checkley71(8403)

SLOPE=(Y2-Y1)/(X2-X1)
SLOPE=(2+3)/(-5+5)
SLOPE=5/0 OR AN UNDEFINED SLOPE WHICH IS A VERTICLE LINE THROUGH X=-5